# FedRecon: Missing Modality Reconstruction in Distributed Heterogeneous Environments

### v1

@@ Abstract.
Multimodal data are often incomplete and exhibit Non-Independent and Identically Distributed (Non-IID) characteristics in real-world scenarios. These inherent limitations lead to both modality heterogeneity through partial modality absence and data heterogeneity from distribution divergence, creating fundamental challenges for effective federated learning (FL). To address these coupled challenges, we propose FedRecon, the first method targeting simultaneous missing modality reconstruction and Non-IID adaptation in multimodal FL. Our approach first employs a lightweight Multimodal Variational Autoencoder (MVAE) to reconstruct missing modalities while preserving cross-modal consistency. Distinct from conventional imputation methods, we achieve sample-level alignment through a novel distribution mapping mechanism that guarantees both data consistency and completeness. Additionally, we introduce a strategy employing global generator freezing to prevent catastrophic forgetting, which in turn mitigates Non-IID fluctuations. Extensive evaluations on multimodal datasets demonstrate FedRecon’s superior performance in modality reconstruction under Non-IID conditions, surpassing state-of-the-art methods.
@@ 1. Introduction
Federated Learning (FL) has gained significant attention as a promising framework for decentralized machine learning, ensuring privacy protection in distributed environments (Li et al., 2020a, c; Qin et al., 2023; Xiong et al., 2024). While existing FL methods demonstrate strong performance in unimodal settings (Chen et al., 2025; Yu et al., 2024b), their inability to capture cross-modal interactions fundamentally limits real-world applicability (Huang et al., 2021, 2024a; Lin et al., 2023). To bridge this gap, Multimodal Federated Learning (MFL) has emerged as a critical solution by enabling complementary information fusion across modalities (Zhao et al., 2022; Che et al., 2023), with successful deployments in autonomous driving (Cui et al., 2019), video action capture (Ofli et al., 2013), and multimedia sentiment analysis (Kaur and Kautish, 2019). By explicitly modeling inter-modal dependencies, MFL achieves holistic understanding of complex phenomena (Li et al., 2024; Chen and Zhang, 2024), enabling precise real-world inference under data scarcity.
@@ 2.2. Missing Modality Reconstruction in MFL
Reconstructing missing modalities in MFL ensures robust model performance under heterogeneous data distributions. Xiong et al. propose CACMRN (Xiong et al., 2023), a client-adaptive transformer leveraging instance relationships and federated optimization to prevent local overfitting during cross-modal reconstruction. Wang et al. develop FedMMR (Wang et al., 2024), aligning feature spaces via modality synthesis while integrating reconstructed and existing modalities as complementary inputs. Liu et al. design FedMobile (Liu et al., 2025), prioritizing cross-node latent feature sharing to address extreme modality incompleteness. Yu et al. introduce FedInMM (Yu et al., 2024a), utilizing LSTM-based dynamic weighting and temporal fusion for irregular multimodal data. While effective in reconstructing modalities under limited samples in MFL, these methods overlook sample-level alignment across clients with identical labels. Additionally, they inadequately address instability caused by Non-IID client data during cross-modal generation. FedRecon stabilizes training by freezing a global generator replica to reduce distributional shifts, thereby mitigating fluctuations induced by client-specific data divergence.
@@ 5.1. Experimental Setups
Implementations. For our MFL framework, we implement FedRecon atop the FedMultimodal benchmark (Feng et al., 2023). Following their protocol, we utilize MobileNetV2 (Howard et al., 2017), MobileBERT (Sun et al., 2020), and MFCC (Mel-Frequency Cepstral Coefficients) (Davis and Mermelstein, 1980) to extract features from visual, textual, and audio modalities respectively, enabling efficient and scalable federated training. Therefore, a simple Conv+RNN architecture is used to process visual and audio data, while text is handled by an RNN-only model. For the MELD dataset, which contains only audio and text modalities, we adapt our model structure based on the VAE design provided in a prior implementation [^1] https://github.com/yjlolo/vae-audio to better handle the audio modality. For CrisisMMD, we adopt the vision-language VAE architecture proposed in the original MVAE framework to accommodate the image-text modality pair. We use FedAvg as the baseline method for MELD, while for CrisisMMD we adopt FedOpt. We report Unweighted Average Recall (UAR) on MELD and F1-score on CrisisMMD, along with Top-1 accuracy on both datasets for a more comprehensive evaluation. To evaluate generative baselines, we build upon the official MMVAE+ implementation [^2] https://github.com/epalu/mmvaeplus and refer to their reported metrics for consistency. We also reproduce MMVM [^3] https://github.com/thomassutter/mmvmvae results as an additional reference point. Generative quality is assessed using Fréchet Inception Distance (FID) (Heusel et al., 2017) and generative coherence (Shi et al., 2019) metrics. All experiments are repeated with three different random seeds to ensure robustness and are conducted on a single NVIDIA A100 GPU with 80GB memory. Due to time constraints, we leave the integration of PolyMNIST and CUB into the FL setting as future work. Further implementation details, including those for reproducing closed-source baselines, are provided in Appendix C.
@@ 5.3. Ablation Studies
Contribution of Each Module. We first evaluate the impact of two key modules: the cross-modal alignment reconstruction generator (MMR) and the Global Generator Freezing Strategy (GGFS). To assess their contributions to model performance, we conduct experiments where these modules are excluded from FedRecon. The results underscore the critical role of these components in enhancing overall effectiveness. As shown in Figure 8, FedRecon demonstrates a marked improvement over the baseline, highlighting the necessity of both MMR and GGFS for superior performance. Due to space limitations, we omit the results on CrisisMMD with $\alpha=5.0$ and on the MELD dataset. On CrisisMMD with $\alpha=5.0$, the MMR and GGFS modules yield average improvements of 1.47% and 0.88% respectively, across varying missing ratios. For MELD, GGFS shows limited or even negative effects, while the MMR module consistently improves UAR by an average of 2.10%.
@@ Appendix C Reproducibility
Due to time constraints, we were unable to organize and release our code repository prior to submission. However, we commit to open-sourcing it upon the paper’s acceptance. The following sections provide detailed technical specifications to facilitate reproducibility.

### v2

# FedRecon: Missing Modality Reconstruction in Heterogeneous Distributed Environments

@@ Abstract.
Multimodal data are often incomplete and exhibit Non-Independent and Identically Distributed (Non-IID) characteristics in real-world scenarios. These inherent limitations lead to both modality heterogeneity through partial modality absence and data heterogeneity from distribution divergence, creating fundamental challenges for effective federated learning (FL). To address these coupled challenges, we propose FedRecon, the first method targeting simultaneous missing modality reconstruction and Non-IID adaptation in multimodal FL. Our approach first employs a lightweight Multimodal Variational Autoencoder (MVAE) to reconstruct missing modalities while preserving cross-modal consistency. Distinct from conventional imputation methods, we achieve sample-level alignment through a novel distribution mapping mechanism that guarantees both data consistency and completeness. Additionally, we introduce a strategy employing global generator freezing to prevent catastrophic forgetting, which in turn mitigates Non-IID fluctuations. Extensive evaluations on multimodal datasets demonstrate FedRecon’s superior performance in modality reconstruction under Non-IID conditions, surpassing state-of-the-art methods. Our code has been released on https://github.com/Wings-Of-Disaster/FedRecon.
@@ 1. Introduction
Federated Learning (FL) has gained significant attention as a promising framework for decentralized machine learning, ensuring privacy protection in distributed environments (Li et al., 2020a, c; Qin et al., 2023; Xiong et al., 2024; Liu et al., 2025a). While existing FL methods demonstrate strong performance in unimodal settings (Chen et al., 2025; Yu et al., 2024b), their inability to capture cross-modal interactions fundamentally limits real-world applicability (Huang et al., 2021, 2024a; Lin et al., 2023). To bridge this gap, Multimodal Federated Learning (MFL) has emerged as a critical solution by enabling complementary information fusion across modalities (Zhao et al., 2022; Che et al., 2023), with successful deployments in autonomous driving (Cui et al., 2019), video action capture (Ofli et al., 2013), and multimedia sentiment analysis (Kaur and Kautish, 2019; Liu et al., 2025b). By explicitly modeling inter-modal dependencies, MFL achieves holistic understanding of complex phenomena (Li et al., 2024; Chen and Zhang, 2024), enabling precise real-world inference under data scarcity.
@@ 2.2. Missing Modality Reconstruction in MFL
Reconstructing missing modalities in MFL ensures robust model performance under heterogeneous data distributions. Xiong et al. propose CACMRN (Xiong et al., 2023), a client-adaptive transformer leveraging instance relationships and federated optimization to prevent local overfitting during cross-modal reconstruction. Wang et al. develop FedMMR (Wang et al., 2024), aligning feature spaces via modality synthesis while integrating reconstructed and existing modalities as complementary inputs. Liu et al. design FedMobile (Liu et al., 2025c), prioritizing cross-node latent feature sharing to address extreme modality incompleteness. Yu et al. introduce FedInMM (Yu et al., 2024a), utilizing LSTM-based dynamic weighting and temporal fusion for irregular multimodal data. While effective in reconstructing modalities under limited samples in MFL, these methods overlook sample-level alignment across clients with identical labels. Additionally, they inadequately address instability caused by Non-IID client data during cross-modal generation. FedRecon stabilizes training by freezing a global generator replica to reduce distributional shifts, thereby mitigating fluctuations induced by client-specific data divergence.
@@ 5.1. Experimental Setups
Implementations. For our MFL framework, we implement FedRecon atop the FedMultimodal benchmark (Feng et al., 2023). Following their protocol, we utilize MobileNetV2 (Howard et al., 2017), MobileBERT (Sun et al., 2020), and MFCC (Mel-Frequency Cepstral Coefficients) (Davis and Mermelstein, 1980) to extract features from visual, textual, and audio modalities respectively, enabling efficient and scalable federated training. Therefore, a simple Conv+RNN architecture is used to process visual and audio data, while text is handled by an RNN-only model. For the MELD dataset, which contains only audio and text modalities, we adapt our model structure based on the VAE design provided in a prior implementation [^1] https://github.com/yjlolo/vae-audio to better handle the audio modality. For CrisisMMD, we adopt the vision-language VAE architecture proposed in the original MVAE framework to accommodate the image-text modality pair. We use FedAvg as the baseline method for MELD, while for CrisisMMD we adopt FedOpt. We report Unweighted Average Recall (UAR) on MELD and F1-score on CrisisMMD, along with Top-1 accuracy on both datasets for a more comprehensive evaluation. To evaluate generative baselines, we build upon the official MMVAE+ implementation [^2] https://github.com/epalu/mmvaeplus and refer to their reported metrics for consistency. We also reproduce MMVM [^3] https://github.com/thomassutter/mmvmvae results as an additional reference point. Generative quality is assessed using Fréchet Inception Distance (FID) (Heusel et al., 2017) and generative coherence (Shi et al., 2019) metrics. All experiments are repeated with three different random seeds to ensure robustness and are conducted on a single NVIDIA A100 GPU with 80GB memory. Further implementation details, including those for reproducing closed-source baselines, are provided in Appendix C.
@@ 5.3. Ablation Studies
Contribution of Each Module. We first evaluate the impact of two key modules: the cross-modal alignment reconstruction generator (MMR) and the Global Generator Freezing Strategy (GGFS). To assess their contributions to model performance, we conduct experiments where these modules are excluded from FedRecon. The results underscore the critical role of these components in enhancing overall effectiveness. As shown in Figure 8, FedRecon demonstrates a marked improvement over the baseline, highlighting the necessity of both MMR and GGFS for superior performance. Owing to space limitations, we exclude detailed results on CrisisMMD with $\alpha=5.0$ and the MELD dataset. Nonetheless, we briefly summarize the findings here. On CrisisMMD ($\alpha=5.0$), MMR and GGFS achieve average performance gains of 1.47% and 0.88% respectively across varying missing rates. For MELD, GGFS shows limited or even negative effects, while the MMR module consistently improves UAR by an average of 2.10%.